# Not all Chess960 positions are equally complex

### v2

@@ Abstract
We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions. Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the initiative is a robust structural feature. To quantify decision difficulty, we introduce an information-based measure $S(n)$ describing the cumulative information required to identify optimal moves over the first $n$ plies. This measure decomposes into contributions from White and Black, $S_{W}$ and $S_{B}$, yielding a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$. Across the ensemble, $S_{\mathrm{tot}}$ varies by a factor of three, while $A$ spans from 2.5-2.5 to +1.8+1.8 bits, showing that some openings burden White and others Black. The mean $\langle A\rangle=-0.25$ bits indicates a slight tendency for White to face harder opening decisions. Standard chess (position #518, RNBQKBNR) exhibits above-average asymmetry (91st percentile) but typical overall complexity (47th percentile). The most complex opening is #226 (BNRQKBNR), whereas #198 (QNBRKBNR)is the most balanced, with both evaluation and asymmetry near zero. These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of the back-rank pieces can significantly alter strategic depth and competitive fairness. Remarkably, the classical starting position—despite centuries of cultural selection—lies far from the most balanced configuration.
@@ I Introduction
Chess, one of humanity’s oldest and most studied strategic games, served as a simple testing ground for theories of decision-making, artificial intelligence, and complex systems [1]. From a physics perspective, chess represents an ideal model system for studying decision-making complexity and information dynamics: it is fully deterministic, has well-defined rules, generates vast empirical datasets, and admits precise computational analysis of optimal strategies. The modern rules of chess crystallized in the 15th century, with the classical starting position—RNBQKBNR (Rook-Knight-Bishop-Queen-King-Bishop-Knight-Rook)—becoming the universal standard. This position was not derived from mathematical principles but emerged through cultural evolution and practical play over centuries [2].
@@
The dominance of opening theory in modern chess has led to a paradoxical situation: at the highest levels of play, the first 15–20 moves often follow extensively analyzed variations stored in databases containing millions of games. This ‘opening book’ knowledge can extend so deeply that the creative, analytical phase of the game is delayed until the middlegame. The tension between memorization and genuine understanding has long concerned chess theorists. As early as 1792, the Dutch chess enthusiast Philip Julius van Zuylen van Nijevelt proposed randomizing starting positions to restore creative play [3]. Two centuries later, World Champion Bobby Fischer recognized the same issue as a potential threat to the game’s intellectual vitality [4]. In 1996, Fischer proposed a variant that preserves chess’s strategic depth while eliminating memorization advantages: Fischer Random Chess, developed in collaboration with Susan Polgar [4]. Now standardized as Chess960 or ‘Fischer random’, and more recently promoted as ‘Freestyle Chess’ [5], this variant shuffles the pieces on the back rank subject to three constraints: (i) the two bishops must occupy opposite-colored squares, (ii) the king must be positioned between the two rooks to preserve castling rights, and (iii) White and Black must share identical mirrored arrangements. These rules generate exactly $960$ legal starting positions, with the classical arrangement RNBQKBNR corresponding to index #518 in the standard numbering.
The number $960$ follows directly from elementary combinatorics. Labeling the back-rank squares (from a to h), the bishops can be placed on any pair of light and dark squares, giving $4\times 4=16$ combinations. The queen may then occupy any of the remaining six squares. This leaves five squares to be filled by the multiset $\{R,K,R,N,N\}$. With 5 squares left, there are $\binom{5}{2}=10$ ways to place the two knights; the remaining three squares must then be occupied by the two rooks and the king, with the king between the rooks, which yields a unique arrangement for these three pieces. Altogether, this gives $16\times 6\times 10=960$ possible starting positions. This deterministic construction also provides a canonical indexing scheme from 0 to $959$, known as the Scharnagl enumeration algorithm [6, 7]. In this scheme, each starting position is uniquely encoded by three integers: (i) the bishop code, selecting one of the $16$ opposite-colored bishop pairs; (ii) the queen position, selecting one of the $6$ remaining squares after the bishops are placed; and (iii) the N5N code, indexing one of the admissible arrangements of the multiset $\{R,K,R,N,N\}$ on the final five squares under the requirement that the king lies between the two rooks. The mapping to an index is given by $\mathrm{idn}\;=\;(\text{bishop code})\;+\;16\times(\text{queen position})\;+\;96\times(\text{N5N code})$. For the classical starting position RNBQKBNR, this yields $\mathrm{idn}=6\;+\;16\times 2\;+\;96\times 5=518$ which is the standard numbering for classical chess. This indexing provides a compact and reproducible method for scanning all $960$ starting configurations and guarantees a one-to-one correspondence between integers in $[0,959]$ and legal Chess960 starting positions.
@@
The variant gained official recognition when FIDE incorporated it into the Laws of Chess in 2008 and began sanctioning world championships in 2019 [4]. More recently, freestyle chess gained more attention with the new ‘Freestyle Chess Grand Slam Tour’ [5]. From the standpoint of statistical physics, Chess960 offers a unique opportunity: a discrete ensemble of 960 initial conditions evolving under identical dynamical rules, enabling systematic comparison of complexity across configurations. While Fischer’s motivation was practical—restoring the importance of understanding over memorization—his innovation raises questions amenable to quantitative analysis: Are all 960 positions equally complex? Does the classical position occupy a special place in the complexity landscape? Can we quantify the ‘decision-making difficulty’ intrinsic to each starting configuration? These questions parallel fundamental inquiries in statistical physics concerning how initial conditions influence the complexity of dynamical trajectories.
@@
Recent advances in computer chess, particularly the development of engines like Stockfish capable of evaluating positions with superhuman accuracy [8], enable us to address these questions rigorously. The availability of large databases has enabled the emergence of a quantitative ‘science of chess,’ with studies addressing innovation dynamics [9], statistical regularities in opening frequencies [10], network properties of move sequences [11, 12, 13, 14, 15], memory effects and temporal correlations [16, 17], game complexity measures [18, 19, 20], human performance and cognition [21, 22, 23], and spatial control patterns [24, 25]. These works demonstrate that chess provides a rich empirical system for studying decision-making under uncertainty.
@@ II Results
II.1 Simple measures
II.1.1 Location probability
@@
As discussed above, there are 960 possibilities that result from the following constraints: The two bishops must occupy opposite-colored squares (one on a light square, one on a dark square); the king must be positioned between the two rooks, preserving the ability to castle kingside and queenside. Finally, we note that the Black pieces mirror the White arrangement, ensuring a priori fairness. These 960 positions are numbered from 0 to 959 and the standard chess position (RNBQKBNR) corresponds to the number 518 in this numbering [26]. The simplest question we can ask under these constraints is: what is the probability that a given piece occupies a particular file? We show this probability in Fig. 1. The bishops, knights, and the queen have equal probability across all eight squares ($1/8=12.5\%$). Due to castling constraints, the king and the rooks have non-uniform probability distributions across the eight squares. The king cannot be on the a or h file and we observe that it is most likely close to the central files (d-e), and naturally, the rooks have a larger probability to be on the sides. Fig. 1 shows the probability distribution for each piece type across the eight squares, revealing the symmetry and constraints of the Chess960 rules.
II.1.2 Initial Stockfish evaluation
@@
To quantify the structural asymmetry inherent in Chess960’s turn-based mechanics, we evaluated all 960 starting positions using Stockfish 17.1 [8] at depth 30. The resulting distribution (Fig. 2) reveals a striking uniformity: the mean initial evaluation is $\langle E\rangle=+0.297\pm 0.136$ pawns, with 956 of 960 positions (99.6%) exhibiting positive evaluation favoring White. Only four position achieve exact or close to balance. This near-universal first-move advantage confirms that the privilege of moving first confers a measurable strategic benefit independent of piece configuration.
@@
Standard chess (position #518, RNBQKBNR) exhibits an initial evaluation of +0.30+0.30 pawns, placing it in the 47th percentile. This demonstrates that the classical starting position is statistically typical within the Chess960 landscape, neither amplifying nor diminishing the structural benefit of initiative.
@@
Figure 2: Distribution of Stockfish initial position evaluations across all 960 Chess960 starting positions (Stockfish 17.1, depth 30). The distribution is centered at $\langle E\rangle=+0.30\pm 0.14$ pawns, with 99.6% of positions favoring White. Vertical lines indicate position #518 (standard chess, solid green), position #279 (maximum White advantage, red dashed), position #535 (almost perfect balance, blue dashed), and the ensemble mean (black dashed).
The distribution’s relatively narrow standard deviation ($\sigma=0.14$ pawns) indicates remarkable consistency across disparate piece arrangements, suggesting that White’s first-move advantage is a robust structural feature largely independent of the specific configuration. The most White-favorable position (#279, NRBKNRQB) reaches +0.83+0.83 pawns—approximately 2.8 times the mean. This configuration places in particular the rook, the queen and a bishop on the f-h files, creating immediate developmental asymmetries that White can exploit (although the engine’s sequence of best moves starting with f4, c4, Nf3 leads to variations that are not obvious). Notably, position #535 (RNBKQNRB) achieves almost perfect balance (with an infinitesimal advantage for Black at 0.09-0.09 pawns) —one of only three configurations where the first-move advantage is close to null. The rarity of balanced or Black-favoring configurations (0.4% of positions) underscores that neutralizing White’s first-move advantage requires a precise alignment of spatial and tactical factors. Note that these results depend on the engine and search depth. While specific outliers may shift, the overall distribution and its statistics remain robust.
@@
II.1.3 Branching factor and position space growth analysis
$b(n)=\frac{N(n+1)}{N(n)}$ (1)
This ratio quantifies the average number of accessible successor states per position at depth $n$. Figure 3 presents a comparative analysis between position 518 (standard chess) and the ensemble average across all sampled positions. Panel (a) shows the evolution of the branching factor $b(n)$ versus ply depth. The average branching factor across sampled positions (blue curve with error bars representing one standard deviation) exhibits typical values of order $20$ consistent with previous estimates [24]. Note that branching factors around $b\sim 35$, often cited in computer chess literature, correspond to middlegame positions where piece development leads to increased move options [24]. Position #518 (green curve) closely tracks this average, with deviations remaining within one standard deviation throughout the analyzed range.
II.2 Characterizing the opening complexity
II.2.1 Asymmetry in Decision Complexity
@@
Consider a position where a chess engine evaluates (in centipawns) the best move with score $E_{1}$ and the second-best move with score $E_{2}$. The difference $\Delta=E_{1}-E_{2}$ quantifies how much ‘better’ the optimal move is [24, 25]. For a player with discrimination ability $\Delta_{0}$ (the minimum evaluation difference they can reliably perceive), we model the probability of choosing the optimal move (over the second best move) using a Boltzmann-like factor (also called the softmax function)
@@
For a trajectory of $n$ plies, the cumulative information required to navigate the resulting branch of the search tree is
@@
with $\Delta_{i}$ the evaluation gap at ply $i$. The average information per move, $S(n)/n$, reflects the typical decision difficulty encountered along the trajectory. $S(n)$ provides a principled, information-theoretic measure of how demanding a given sequence of moves (such as an opening for example) is to navigate under optimal play.
The quantity $S(n)$ has several natural properties: it is monotonically increasing with $n$ (information accumulates along the line of play); forced positions with large $\Delta_{i}$ contribute negligibly; subtle positions with small $\Delta_{i}$ dominate; and the overall magnitude depends on playing strength through $\Delta_{0}$. Typical values are $\Delta_{0}\approx 10$ cp for Grandmaster level, $\Delta_{0}\approx 50$ cp for Expert level, and $\Delta_{0}\approx 100$ cp for intermediate level [25]. We focus on the Grandmaster level here and set $\Delta_{0}=10$ cp.
Figure 4: Information cost analysis across the 960 Chess960 starting positions based on $n=10$ plies per player of optimal play (Stockfish 17.1, depth 14, discrimination threshold $\Delta_{0}=10$ cp). Values represent averages over 5 games per position. (a) Distribution of White’s cumulative cost $S_{W}(10)$ showing the decision complexity faced by White across different starting configurations. The green vertical line marks position #518 (standard chess) at $\langle S_{W}\rangle=4.29$ bits, near the ensemble mean. (b) Distribution of Black’s cumulative cost $S_{B}(10)$. Position #518 exhibits $\langle S_{B}\rangle=4.97$ bits, slightly above the ensemble mean. (c) Asymmetry $A=S_{B}-S_{W}$ versus total information cost $S_{\mathrm{tot}}=S_{W}+S_{B}$. Negative asymmetry indicates White faces more complex decisions; positive values indicate Black faces greater complexity. Standard chess (position #518, indicated by a gold star) shows positive asymmetry ($\langle A\rangle=+0.69$ bits) and moderate total complexity (46.5th percentile). (d) Distribution of asymmetry across all positions. The mean asymmetry is $\langle A\rangle=-0.25$ bits (and standard deviation $1.42$ bits), suggesting a weak structural advantage for Black across Chess960 configurations. Standard chess lies at the 91.2th percentile for asymmetry.
@@
The information cost framework reveals significant variation in decision-making burden between White and Black across Chess960 positions. We define the asymmetry as

### v3

@@ Abstract
We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions. Stockfish evaluations reveal a near-universal first-move advantage for White ($\langle E\rangle=+0.33\pm 0.12$ pawns), indicating that the initiative is a robust structural feature of the game. To quantify decision difficulty, we introduce an information-based measure $S(n)$ that captures the cumulative information required to identify optimal moves over the first $n$ plies. This measure decomposes into White and Black contributions, $S_{W}$ and $S_{B}$, defining a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$. Across the ensemble, $S_{\mathrm{tot}}$ ranges from $2.6$ to $17.2$ bits, while $A$ spans from 4.5-4.5 to +4.2+4.2 bits (mean $\langle A\rangle=-0.26$), showing that openings are nearly evenly split between those that burden White and those that burden Black, with a slight average excess complexity for White. Standard chess (position #518, RNBQKBNR) exhibits near-average total complexity and asymmetry, yet lies far from the configuration that jointly minimizes evaluation imbalance and decision asymmetry. These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of back-rank pieces can substantially alter strategic depth and competitive balance. The classical starting position—despite centuries of refinement—appears not as an extremum, but as one configuration among many in a broad statistical ensemble.
@@ I Introduction
Chess, one of humanity’s oldest and most studied strategic games, served as a simple testing ground for theories of decision-making, artificial intelligence, and complex systems Levy (1988). From a physics perspective, chess represents an ideal model system for studying decision-making complexity and information dynamics: it is fully deterministic, has well-defined rules, generates vast empirical datasets, and admits precise computational analysis of optimal strategies. The modern rules of chess crystallized in the 15th century, with the classical starting position—RNBQKBNR (Rook-Knight-Bishop-Queen-King-Bishop-Knight-Rook)—becoming the universal standard. This position was not derived from mathematical principles but emerged through cultural evolution and practical play over centuries Murray (1917).
@@
The dominance of opening theory in modern chess has led to a paradoxical situation: at the highest levels of play, the first 15–20 moves often follow extensively analyzed variations stored in databases containing millions of games. This ‘opening book’ knowledge can extend so deeply that the creative, analytical phase of the game is delayed until the middlegame. The tension between memorization and genuine understanding has long concerned chess theorists. As early as 1792, the Dutch chess enthusiast Philip Julius van Zuylen van Nijevelt proposed randomizing starting positions to restore creative play van Zuylen van Nijevelt (1792). Two centuries later, World Champion Bobby Fischer recognized the same issue as a potential threat to the game’s intellectual vitality Polgar (2025). In 1996, Fischer proposed a variant that preserves chess’s strategic depth while eliminating memorization advantages: Fischer Random Chess, developed in collaboration with Susan Polgar Polgar (2025). Now standardized as Chess960 or ‘Fischer random’, and more recently promoted as ‘Freestyle Chess’ 13, this variant shuffles the pieces on the back rank subject to three constraints: (i) the two bishops must occupy opposite-colored squares, (ii) the king must be positioned between the two rooks to preserve castling rights, and (iii) White and Black must share identical mirrored arrangements. These rules generate exactly $960$ legal starting positions (see Appendix A), with the classical arrangement RNBQKBNR corresponding to index #518 in the standard numbering.
@@
Chess960 gained official recognition when FIDE incorporated it into the Laws of Chess in 2008 and began sanctioning world championships in 2019 Polgar (2025). More recently, it received even more attention with the new ‘Freestyle Chess Grand Slam Tour’ 13. From the standpoint of statistical physics, Chess960 offers a unique opportunity: a discrete ensemble of 960 initial conditions evolving under identical dynamical rules, enabling systematic comparison of complexity across configurations. While Fischer’s motivation was practical—restoring the importance of understanding over memorization—his innovation raises questions amenable to quantitative analysis: Are all 960 positions equally complex? Does the classical position occupy a special place in the complexity landscape? Can we quantify the ‘decision-making difficulty’ intrinsic to each starting configuration? These questions parallel fundamental inquiries in statistical physics concerning how initial conditions influence the complexity of dynamical trajectories.
@@
Recent advances in computer chess, particularly the development of engines like Stockfish capable of evaluating positions with superhuman accuracy Stockfish Development Team (2024), enable us to address these questions rigorously. The availability of large databases has enabled the emergence of a quantitative ‘science of chess,’ with studies addressing innovation dynamics Perotti et al. (2013), statistical regularities in opening frequencies Blasius and Tonjes (2009), network properties of move sequences Maslov (2009); Farren et al. (2013); Ribeiro et al. (2013); Barthelemy (2025b); Cerioli et al. (2025), memory effects and temporal correlations Schaigorodsky et al. (2014, 2016), game complexity measures Atashpendar et al. (2016); De Marzo and Servedio (2023); Chacoma and Billoni (2024), human performance and cognition Chowdhary et al. (2023); Chassy and Gobet (2011); Sigman et al. (2010), and spatial control patterns Barthelemy (2023, 2025a). These works demonstrate that chess provides a rich empirical system for studying decision-making under uncertainty.
@@ II Results
In this section, we present three complementary indicators characterizing the 960 starting positions: (i) the positional probability distribution of pieces in the initial configuration, (ii) the engine-based structural evaluation of all positions, and (iii) a measure of the complexity burden borne by White and Black during the first moves of the game.
While many other metrics could in principle be considered, our objective here is to focus on quantities directly related to structural imbalance and early-game decision difficulty. Additional measures—such as the branching factor—are briefly discussed in Appendix B, where we explain why they are less central to the questions addressed in this study.
II.1 Location probability
@@
The 960 possibilities result from the following constraints (see Appendix A): The two bishops must occupy opposite-colored squares (one on a light square, one on a dark square); the king must be positioned between the two rooks, preserving the ability to castle kingside and queenside. Finally, Black pieces mirror the White arrangement, ensuring a priori fairness. These 960 positions are numbered from 0 to 959 and the standard chess position (RNBQKBNR) corresponds to the number 518 in this numbering 9. The simplest question we can ask under these constraints is: what is the probability that a given piece occupies a particular file? We show this probability in Fig. 1. The bishops, knights, and the queen have equal probability across all eight squares ($1/8=12.5\%$). Due to castling constraints, the king and the rooks have non-uniform probability distributions across the eight squares. The king cannot be on the a or h file and we observe that it is most likely close to the central files (d-e), and naturally, the rooks have a larger probability to be on the sides. Fig. 1 shows the probability distribution for each piece type across the eight squares, revealing the symmetry and constraints of the Chess960 rules.
II.2 Initial Stockfish evaluation
In order to quantify the intrinsic structural asymmetry of Chess960 starting positions, we perform a systematic engine evaluation of all 960 configurations. The goal is not to model human perception or practical playing strength, but to estimate the underlying structural advantage encoded in each initial arrangement under near-optimal continuation. To this end, we evaluated all 960 starting positions using Stockfish 17.1 Stockfish Development Team (2024) at fixed depth 30. The depth-30 score is interpreted here as an engine-level proxy for structural advantage, reflecting the evaluation after a sufficiently deep search while remaining computationally tractable. All computations were performed with NNUE enabled, using a single search thread (Threads=1) and a 1,GB hash table (Hash=1024). Each position was analyzed under a strict depth constraint without time limits, ensuring that every evaluation reached the prescribed search depth exactly.
@@ II.2 Initial Stockfish evaluation
The resulting distribution (Fig. 2) reveals a striking uniformity: the mean initial evaluation is $\langle E\rangle=+0.33\pm 0.12$ pawns, with 959 of 960 positions (99.9%) exhibiting positive evaluation favoring White. Only one position (#774) achieves a small avantage for black (E=0.18E=-0.18). This near-universal first-move advantage confirms that the privilege of moving first confers a measurable strategic benefit independent of piece configuration.
@@
Standard chess (position #518, RNBQKBNR) exhibits an initial evaluation of +0.28+0.28 pawns, placing it in the 37th percentile. This demonstrates that the classical starting position is statistically typical within the Chess960 landscape, neither amplifying nor diminishing the structural benefit of initiative.
@@
Figure 2: Distribution of Stockfish initial evaluations across all 960 Chess960 starting positions (Stockfish 17.1, depth 30, Threads=1). The distribution is centered at $\langle E\rangle=+0.33\pm 0.12$ pawns, with $99.9\%$ of positions favoring White. Vertical lines indicate position #518 (standard chess, solid green). At this depth, position #333 exhibits the maximum White advantage (red dashed), while position #774 is the closest to perfect evaluation balance (blue dashed). The ensemble mean is shown as a black dashed line.
The distribution’s relatively narrow standard deviation ($\sigma=0.12$ pawns) indicates remarkable consistency across disparate piece arrangements, suggesting that White’s first-move advantage is a robust structural feature largely independent of the specific configuration.
The most White-favorable position (#333, NRQBKRBN) reaches the surprisingly large value of +0.81+0.81 pawns—approximately 2.5 times the mean. Inspection of the principal variation reveals that White can rapidly mobilize its pieces with the plan Nb3, f4, and Bc5, seizing central and kingside space while targeting the weakened dark squares around Black’s king. The structural origin of this asymmetry might lie in the initial placement itself: White’s queen on c1 and bishop on d1 are already aligned along natural attacking diagonals, while the knight on a1 finds an efficient route to b3 where it supports both the center and the c5 outpost. More analysis of the variations are however necessary to inspect in more details this asymmetry. Another interesting case is the only position #774 (QRBKNBRN) that gives an advantage to black, although very small (0.18-0.18). The rarity of balanced configurations underscores that neutralizing White’s first-move advantage requires a precise alignment of spatial and tactical factors.
@@
II.3 Effect of depth: instability of extremes
The outliers identified in the previous section were obtained at depth $30$. A natural question is whether these extreme evaluations remain stable as the search depth varies. In Fig. 3, we plot the three positions with the largest evaluations and the three with the smallest evaluations (as previously computed with Stockfish 17.1, Threads=1, NNUE enabled, identical hardware and hash settings) as a function of search depth.
Figure 3: (Top) Evaluation of the three highest- and three lowest-ranked initial positions as a function of search depth. (Bottom) Evaluation of position #333 as a function of depth. All evaluations are expressed in pawns from White’s perspective.
The result is clear: extreme positions are not depth-stable. While aggregate statistics (mean, variance, overall distribution shape) remain robust, individual outliers fluctuate substantially, as illustrated in Fig. 3(top). At shallow depth the instability is striking. At depth 10, the top configuration (position 702) reaches an evaluation of +5.89+5.89 pawns, followed by +3.41+3.41 and +2.19+2.19. These extreme values completely disappear at greater depth. Already at depth 15, the top evaluations drop to $\approx+1.16$, and by depth 30 the maximum is +0.81+0.81. At depth 40, the leading positions cluster tightly around +0.79+0.79. Thus, the apparent large advantages observed at depth 10 are largely artifacts of shallow search. For bottom positions variations are much smaller. At depth 10, the lowest evaluations are only mildly negative (0.21-0.21, 0.16-0.16, 0.10-0.10), but the identities of these configurations change repeatedly with depth. For instance, position 774 appears among the bottom three at depths 30 and 40, while other initially weak positions disappear from the extreme tail. Moreover, several configurations that are slightly negative at shallow depth become approximately neutral at intermediate depths, before stabilizing again.
The turnover in the ranking is equally informative. None of the top three positions at depth 10 remains in the top three at depth 20 or beyond. Conversely, position #333—identified as the strongest configuration at depth 30 (+0.81+0.81)—does not appear among the leaders at shallow depth. Even between depths 30 and 40, although the numerical evaluations become closer, the composition of the top group still changes. To illustrate this sensitivity, we focus specifically on position #333 and display in Fig. 3 its evaluation as a function of search depth. The figure confirms that the evaluation is not stable with depth and does not exhibit clear convergence within the explored range. While the advantage peaks around depths 25–30, it subsequently decreases, reaching approximately +0.5+0.5 pawns at depth 40.
This behavior indicates that extreme evaluations are particularly sensitive to deeper tactical or strategic resources revealed by extended search. Positions that appear highly favorable at shallow depth may rely on lines that are later mitigated by longer forcing sequences or defensive resources uncovered at greater depth. More generally, this raises two questions: (i) do some positions converge rapidly and remain stable across depths, while others exhibit strong depth sensitivity? and (ii) which structural features control this stability? Addressing these issues would require a systematic study and the introduction of a quantitative measure of evaluation stability, allowing positions to be classified according to their sensitivity to search-horizon effects.
In summary, while global evaluation statistics are robust, the ranking of extreme Chess960 starting positions is not invariant with respect to search depth. Claims about the “most advantageous” or “most disadvantageous” configurations should therefore be supported by an explicit depth-stability analysis. Search depth should be interpreted as defining the resolution at which the decision landscape is probed.
II.4 Characterizing the opening complexity
@@
Consider a position where a chess engine evaluates (in centipawns) the best move with score $E_{1}$ and the second-best move with score $E_{2}$. The difference $\Delta=E_{1}-E_{2}$ quantifies how much ‘better’ the optimal move is Barthelemy (2023, 2025a). For a player with discrimination ability $\Delta_{0}$ (the minimum evaluation difference they can reliably perceive), we model the probability $P(\text{optimal})$ of choosing the optimal move (over the second best move) using a Boltzmann-like factor (also called the softmax function)
@@
For a trajectory of $n$ plies, the cumulative information required to navigate the corresponding branch of the search tree is
@@
where $\Delta_{i}$ denotes the evaluation gap at ply $i$. The average information per move, $S(n)/n$, characterizes the typical decision difficulty encountered along the trajectory. Thus, $S(n)$ provides an information-theoretic measure of how demanding it is to follow a given sequence of moves (for example, an opening line) under optimal play. The quantity $S(n)$ has several natural properties. It is monotonically increasing with $n$, reflecting the cumulative nature of information along the line of play. Positions with large $\Delta_{i}$ (quasi-forced moves) contribute negligibly, whereas positions with small $\Delta_{i}$ dominate the sum. The overall scale of $S(n)$ depends on the discrimination parameter $\Delta_{0}$, which encodes the sensitivity to small evaluation differences. In Appendix C, we further show that $S(10)$ correlates with the mean thinking time $\tau$ of human players: $\tau$ increases with $S(10)$, supporting its interpretation as a meaningful proxy for decision-making complexity. Importantly, $S(n)$ is not intended to model human computational limits. Rather, it quantifies the intrinsic difficulty of identifying the optimal path among competing continuations (here in the early game). In this sense, $S(n)$ is a structural property of the position: it measures how sharply the optimal move is separated from alternatives along the principal variation. It is therefore independent of player strength and instead characterizes the geometry of the decision landscape itself.
Estimating a precise value for $\Delta_{0}$ is non-trivial, and any choice is necessarily somewhat arbitrary. In Appendix C, we show that the percentile ranking of a representative position (the standard position #518) varies only weakly with $\Delta_{0}$, indicating robustness with respect to this parameter. We also provide empirical evidence that $\Delta_{0}\approx 10$ cp corresponds to a natural discrimination scale at expert (Grandmaster) level. In what follows, we therefore focus on this high-skill regime and set $\Delta_{0}=10$ cp.
Figure 4: Information cost analysis across the 960 Chess960 starting positions, computed for $n=10$ plies per player under optimal play (Stockfish 17.1, depth 15, discrimination threshold $\Delta_{0}=10$ cp). (a) Distribution of White’s cumulative cost $S_{W}(10)$, quantifying the decision complexity faced by White across starting configurations. The green vertical line marks position #518 (standard chess), with $\langle S_{W}\rangle=5.42$ bits. (b) Distribution of Black’s cumulative cost $S_{B}(10)$. Position #518 yields $\langle S_{B}\rangle=5.79$ bits, slightly above the ensemble mean. (c) Asymmetry $A=S_{B}-S_{W}$ versus total information cost $S_{\mathrm{tot}}=S_{W}+S_{B}$. Negative asymmetry indicates greater decision complexity for White; positive values indicate greater complexity for Black. The average asymmetry is slightly negative, $\langle A\rangle=-0.26$ bits (standard deviation $1.36$), while the average total cost is $\langle S_{\mathrm{tot}}\rangle=9.69$ bits (standard deviation $2.4$). Standard chess (position #518, gold star) exhibits positive asymmetry ($A_{518}=+0.37$ bits) and a relatively high total cost ($11.2$ bits, 72.5th percentile). (d) Distribution of asymmetry across all positions. The mean asymmetry is $\langle A\rangle=-0.26$ bits, indicating a slight average structural complexity advantage for Black across Chess960. Standard chess lies at the 69th percentile of the asymmetry distribution, placing it close to the ensemble average.
@@
The information cost framework reveals important variations in decision-making burden between White and Black across Chess960 positions. We define the asymmetry as